# CTBC: Contact-Triggered Blind Climbing for Wheeled Bipedal Robots with Instruction Learning and Reinforcement Learning

### v2

@@ V CONCLUSIONS
In conclusion, we propose a contact-triggered, blind stair-climbing method for biped-wheeled robots that blends feedforward instruction learning with reinforcement learning, enabling the robot to conquer stairs and holes without any external perception. Both simulation and hardware experiments demonstrate robust traversal over stairs and holes, confirming the framework’s terrain-crossing capability. However, the extended front–rear gait in the first phase limits the rear leg’s contact opportunities, causing the second-phase policy to consistently lift the front leg first and become locked into this early preference. Moreover, the current strategy remains purely blind and has yet to incorporate external perception, navigation, or trajectory planning. Future work will therefore focus on eliminating this early bias and adopting the blind policy as a low-level controller, augmented by a lightweight vision/LiDAR high-level policy. This hierarchical “blind–perceptive” architecture will enable real-time terrain estimation, global navigation, and efficient traversal of stairs and holes in unknown environments.

### v3

@@ V CONCLUSIONS
In conclusion, this paper presents a contact-triggered, blind locomotion framework for bipedal wheeled-legged robots, successfully bridging the gap between prescribed feedforward guidance and adaptive reinforcement learning. By internalizing motion primitives through a two-stage training scheme, the robot achieves robust stair-climbing and hole-traversal without any exteroceptive sensing. Experimental validations on both the 20 kg LimX Dynamics Tron1 and the 65 kg Cowarobot R0 demonstrate the framework’s exceptional cross-hardware versatility. Despite significant discrepancies in physical scale, wheel diameters, and tire materials (pneumatic vs. solid rubber), the policy generalizes effectively to obstacles multiple times the height of the initial guidance.
Nevertheless, certain limitations remain: the extended gait in the first training phase restricts the contact frequency of the rear legs, leading to a persistent kinematic bias where the policy favors leading with the front legs. Furthermore, the current system remains purely blind, limiting its efficiency in complex, unstructured environments. Future work will explore symmetry-breaking rewards to eliminate gait bias and integrate this robust blind policy as a low-level reactive controller within a hierarchical perceptive architecture. This synergy between ”blind-reflex” and ”vision-based planning” will ultimately empower bipedal wheeled-legged robots with fully autonomous navigation and traversal capabilities in unknown and hazardous terrains.